Estimate the probability that a 100-eV neutron will penetrate a 1 cm 10B (Boron-10) shield and produce a fission in a 1-mm 239Pu foil (density 18.5 g cm–3) behind it. Neglect energy loss of the neutron due to elastic scattering.
Answer #1

Given;

Neutron energy=100eV

=1240/100 nm

=12.4*10^-7cm

The penetrated element given is Boron (of mass no.10) being penetrated at a length = 1 cm

This produces a fission in Plutonium foil (of mass no.239) of density = 18.5gcm^-3

Since it penetrated the region in element that causes fission=239/18.5Vdv =6.45

Now Estimating probability;

Neutron energy/ region =12.4*10^-7/6.45

=1.9*10^-7

In percentage; the answer is 1.9*10^-5